### v1

@@ Appendix B Proof of Proposition 1
For any fixed $\mathcal{F},\mathcal{C}\subset[1:n]$ satisfying $|\mathcal{C}|\leq c<t\leq|\mathcal{F}|\leq n$, and for any $\sigma^{2}$, $r$, the optimum decoder is given by,
$\bm{\alpha}^{*}=\arg\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\bm{\alpha})=\frac{1}{1+\frac{d}{|\mathcal{U}|}(\sigma^{2}+r(|\mathcal{F}|-1))}\mathbf{1}_{|\mathcal{U}|}$ (25)
Proof:  Let $\mathcal{U}=\mathcal{F}\setminus\mathcal{C}$, and let $x_{i,k}$ and $Z_{i,k}$ denote the $k$th element of $\mathbf{x}_{i}$ and $\mathbf{Z}_{i}$, respectively. Then,
@@
$\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\bm{\alpha})$ $=\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}\mathbb{E}\left[\left\|\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\alpha_{i}(\mathbf{x}_{i}+\mathbf{Z}_{i})-\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\mathbf{x}_{i}\right\|^{2}\right]$ (26)
$=\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}\sum_{k=1}^{d}\mathbb{E}\left[\left(\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\alpha_{i}({x}_{i,k}+Z_{i,k})-\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}{x}_{i,k}\right)^{2}\right]$ (27)
$=\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}\frac{1}{|\mathcal{U}|^{2}}\sum_{k=1}^{d}\left(\bm{\alpha}^{T}A_{k}\bm{\alpha}-2\cdot\mathbf{1}^{T}\mathbf{x}^{[k]}\mathbf{x}^{[k]T}\bm{\alpha}+\mathbf{1}^{T}\mathbf{x}^{[k]}\mathbf{x}^{[k]T}\mathbf{1}\right)$ (28)
where $\mathbf{x}^{[k]}=[x_{j_{1},k},\dotsc,x_{j_{|\mathcal{U}|},k}]^{T}$ and $A_{k}=\mathbf{x}^{[k]}\mathbf{x}^{[k]T}+\Sigma$, where $\Sigma$ is the covariance matrix of $[Z_{1,k},\dotsc,Z_{|\mathcal{U}|,k}]^{T}$ for all $k\in[1:d]$. Define,
$f(\bm{\alpha},\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}},\Sigma)=\frac{1}{|\mathcal{U}|^{2}}\sum_{k=1}^{d}\left(\bm{\alpha}^{T}A_{k}\bm{\alpha}-2\cdot\mathbf{1}^{T}\mathbf{x}^{[k]}\mathbf{x}^{[k]T}\bm{\alpha}+\mathbf{1}^{T}\mathbf{x}^{[k]}\mathbf{x}^{[k]T}\mathbf{1}\right).$ (29)
For any fixed $\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}}$ and $\Sigma$, $f(\bm{\alpha},\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}},\Sigma)$ is convex in $\bm{\alpha}$ as all $A_{k}$, $k\in[1:d]$ are positive definite. For any given $\bm{\alpha}$ in (26), as $\mathbf{Z}_{i}$, $i\in\mathcal{U}$ are i.i.d. and $\mathbf{x}_{i}$, $i\in\mathcal{U}$ are chosen from the same set $\mathbb{B}^{d}$, we have,
$\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\bm{\alpha})$ $=\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\Pi_{j}(\bm{\alpha})),\quad j\in[1:|\mathcal{U}|!]$ (30)
where $\Pi_{j}(\bm{\alpha})$ denotes the $j$th permutation of the elements of $\bm{\alpha}$. Therefore,
$\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\bm{\alpha})$ $=\frac{1}{|\mathcal{U}|!}\sum_{j=1}^{|\mathcal{U}|!}\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\Pi_{j}(\bm{\alpha}))$ (31)
@@
$=\frac{1}{|\mathcal{U}|!}\sum_{j=1}^{|\mathcal{U}|!}\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}f\left(\Pi_{j}(\bm{\alpha}),\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}},\Sigma\right)$ (32)
@@
$\geq\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}\frac{1}{|\mathcal{U}|!}\sum_{j=1}^{|\mathcal{U}|!}f\left(\Pi_{j}(\bm{\alpha}),\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}},\Sigma\right)$ (33)
@@
$\geq\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}f\left(\frac{1}{|\tau|!}\sum_{j=1}^{|\tau|!}(\Pi_{j}(\bm{\alpha})),\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}},\Sigma\right)$ (34)
$=\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}f\left(\mathbf{\tilde{\alpha}},\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}},\Sigma\right)$ (35)
where (34) is due to the convexity of $f(\bm{\alpha},\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}},\Sigma)$ in $\bm{\alpha}$, and $\mathbf{\tilde{\alpha}}=\left(\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\alpha_{i}\right)\mathbf{1}_{|\mathcal{U}|}$. This implies that for any decoder $\bm{\alpha}$ with $\mathcal{U}$, $\sigma^{2}$ and $r$ fixed, there exists a $\mathbf{\tilde{\alpha}}=[\tilde{\alpha}_{j_{1}},\dotsc,\tilde{\alpha}_{j_{|\mathcal{U}|}}]^{T}$ such that $\tilde{\alpha}_{k}=\tilde{\alpha}_{\ell}$, $\forall k,\ell$ satisfying $\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\bm{\alpha})\geq\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\mathbf{\tilde{\alpha}})$. Therefore, the optimum decoder is of the form $\bm{\alpha}=\alpha\mathbf{1}_{|\mathcal{U}|}$, and (26) can be written as,
$\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\alpha)$
@@
$=\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}\mathbb{E}\left[\left\|\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\alpha(\mathbf{x}_{i}+\mathbf{Z}_{i})-\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\mathbf{x}_{i}\right\|^{2}\right]$ (36)
@@
$=\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}\frac{1}{|\mathcal{U}|^{2}}\sum_{k=1}^{d}\left((\alpha-1)^{2}\sum_{i\in\mathcal{U}}x_{i,k}^{2}+2(\alpha-1)^{2}\sum_{\begin{subarray}{c}i,j\in\mathcal{U}\\ i\neq j\end{subarray}}x_{i,k}x_{j,k}\right.$
$\qquad\qquad\qquad\qquad\qquad\qquad\left.+2\alpha^{2}\sum_{\begin{subarray}{c}i,j\in\mathcal{U}\\ i\neq j\end{subarray}}r+\alpha^{2}(\sigma^{2}+r|\mathcal{F}\cap\mathcal{C}|)|\mathcal{U}|\right)$ (37)
@@
$=\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}\frac{1}{|\mathcal{U}|^{2}}\left((\alpha-1)^{2}\sum_{i\in\mathcal{U}}\|\mathbf{x}_{i}\|^{2}+2(\alpha-1)^{2}\sum_{\begin{subarray}{c}i,j\in\mathcal{U}\\ i\neq j\end{subarray}}\mathbf{x}_{i}^{T}\mathbf{x}_{j}+d\alpha^{2}r|\mathcal{U}|(|\mathcal{U}|-1)\right.$
$\qquad\qquad\qquad\qquad\qquad\qquad\left.+d\alpha^{2}(\sigma^{2}+r|\mathcal{F}\cap\mathcal{C}|)|\mathcal{U}|\right)$ (38)
where (37) comes from $\text{Var}(\mathbf{Z}_{i})=-r(n-1-|\mathcal{F}\cap\mathcal{C}|)+\sigma^{2}+r(n-1)=\sigma^{2}+r|\mathcal{F}\cap\mathcal{C}|$. The worst case MSE is resulted when $\mathbf{x}_{i}^{T}\mathbf{x}_{j}=1$, for all $i,j\in\mathcal{U}$, that is, when $\mathbf{x}_{i}=\mathbf{x}_{j}$, $\forall i,j$, and $\mathbf{x}_{i}\in\mathbb{S}^{d-1}$, where $\mathbb{S}^{d-1}$ is the unit sphere in $\mathbb{R}^{d}$. Therefore,
$\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\alpha)$
@@
$\quad=\frac{1}{|\mathcal{U}|^{2}}\left((\alpha-1)^{2}|\mathcal{U}|+(\alpha-1)^{2}|\mathcal{U}|(|\mathcal{U}|-1)+d\alpha^{2}r|\mathcal{U}|(|\mathcal{U}|-1)+d\alpha^{2}(\sigma^{2}+r|\mathcal{F}\cap\mathcal{C}|)|\mathcal{U}|\right)$ (39)
@@
$\quad=(\alpha-1)^{2}+\frac{d\alpha^{2}}{|\mathcal{U}|}(\sigma^{2}+r(|\mathcal{F}|-1))$ (40)
as $|\mathcal{U}|+|\mathcal{F}\cap\mathcal{C}|=|\mathcal{F}|$. The optimum decoder for any fixed $\mathcal{U}$, $\sigma^{2}$ and $r$ is computed as,
$\frac{\partial\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\alpha)}{\partial\alpha}$ $=2(\alpha-1)+\frac{2d\alpha}{|\mathcal{U}|}(\sigma^{2}+r(|\mathcal{F}|-1))=0$ (41)
$\alpha^{*}$ $=\frac{1}{1+\frac{d}{|\mathcal{U}|}(\sigma^{2}+r(|\mathcal{F}|-1))}.$ (42)

### v2

@@ Appendix B Proof of Proposition 1
Proposition 1 restated: For any fixed $\mathcal{U}\subseteq\mathcal{U}_{all}$ satisfying $|\mathcal{U}|\geq t$, and for any $\sigma^{2}$, $r$, the optimum decoder is given by,
$\boldsymbol{\alpha}^{*}_{\mathcal{U}}$ $=\arg\min_{\boldsymbol{\alpha}_{\mathcal{U}}}\mathrm{MSE}(\sigma^{2},\rho,\mathcal{U},\boldsymbol{\alpha}_{\mathcal{U}})$ (43)
$=\frac{1}{1+\frac{d}{|\mathcal{U}|}(\sigma^{2}+r(|\mathcal{U}|-1))}\mathbf{1}_{|\mathcal{U}|}$ (44)
The corresponding MSE is given by,
@@
$\min_{\boldsymbol{\alpha}_{\mathcal{U}}}\mathrm{MSE}(\sigma^{2},\rho,\mathcal{U},\boldsymbol{\alpha}_{\mathcal{U}})=\left(1+\frac{|\mathcal{U}|/d}{\sigma^{2}(1+\rho(|\mathcal{U}|-1))}\right)^{-1}$ (45)
Proof:  Let $x_{i,k}$ and $Z_{i,k}$ denote the $k$th element of $\mathbf{x}_{i}$ and $\mathbf{Z}_{i}$, respectively. Then,
$\mathrm{MSE}(\sigma^{2},\rho,\mathcal{U},\boldsymbol{\alpha}_{\mathcal{U}})$
@@
$=\sup_{\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}}\in\mathbb{B}^{d}}\mathbb{E}\!\!\left[\left\|\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\alpha_{i}(\mathbf{x}_{i}+\mathbf{Z}_{i})\!-\!\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\mathbf{x}_{i}\right\|^{2}\right]$ (46)
@@
$=\!\!\!\!\!\!\sup_{\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}}\in\mathbb{B}^{d}}\sum_{k=1}^{d}\mathbb{E}\!\!\left[\!\left(\!\frac{1}{|\mathcal{U}|}\!\sum_{i\in\mathcal{U}}\alpha_{i}({x}_{i,k}\!+\!Z_{i,k})\!-\!\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}{x}_{i,k}\!\right)^{2}\!\right]$ (47)
@@
$=\sup_{\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}}\in\mathbb{B}^{d}}\frac{1}{|\mathcal{U}|^{2}}\sum_{k=1}^{d}\left(\boldsymbol{\alpha}_{\mathcal{U}}^{T}A_{k}\boldsymbol{\alpha}_{\mathcal{U}}-2\cdot\mathbf{1}^{T}\mathbf{x}^{[k]}\mathbf{x}^{[k]T}\boldsymbol{\alpha}_{\mathcal{U}}\right.$
$\qquad\qquad\qquad\qquad\qquad\qquad\left.+\mathbf{1}^{T}\mathbf{x}^{[k]}\mathbf{x}^{[k]T}\mathbf{1}\right)$ (48)
where $\mathbf{x}^{[k]}=[x_{j_{1},k},\dotsc,x_{j_{|\mathcal{U}|},k}]^{T}$ and $A_{k}=\mathbf{x}^{[k]}\mathbf{x}^{[k]T}+\Sigma$, where $\Sigma$ is the covariance matrix of $[Z_{j_{1},k},\dotsc,Z_{j_{|\mathcal{U}|},k}]^{T}$ for all $k\in[1:d]$. Define,
$f(\boldsymbol{\alpha}_{\mathcal{U}},\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}},\Sigma)$
@@
$=\frac{1}{|\mathcal{U}|^{2}}\sum_{k=1}^{d}\left(\boldsymbol{\alpha}_{\mathcal{U}}^{T}A_{k}\boldsymbol{\alpha}_{\mathcal{U}}-2\cdot\mathbf{1}^{T}\mathbf{x}^{[k]}\mathbf{x}^{[k]T}\boldsymbol{\alpha}_{\mathcal{U}}+\mathbf{1}^{T}\mathbf{x}^{[k]}\mathbf{x}^{[k]T}\mathbf{1}\right)$ (49)
For any fixed $\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}}$ and $\Sigma$, $f(\boldsymbol{\alpha}_{\mathcal{U}},\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}},\Sigma)$ is convex in $\boldsymbol{\alpha}_{\mathcal{U}}$ as all $A_{k}$, $k\in[1:d]$ are positive definite. For any given $\boldsymbol{\alpha}_{\mathcal{U}}$ in (B), as $\mathbf{Z}_{i}$, $i\in\mathcal{U}$ are i.i.d. and $\mathbf{x}_{i}$, $i\in\mathcal{U}$ are chosen from the same set $\mathbb{B}^{d}$, we have,
$\mathrm{MSE}(\sigma^{2},\rho,\mathcal{U},\boldsymbol{\alpha}_{\mathcal{U}})$ $=\mathrm{MSE}(n,\mathcal{U},\sigma^{2},\rho,\Pi_{j}(\boldsymbol{\alpha}_{\mathcal{U}})),$ (50)
for $j\in[1:|\mathcal{U}|!]$ where $\Pi_{j}(\boldsymbol{\alpha}_{\mathcal{U}})$ denotes the $j$th permutation of the elements of $\boldsymbol{\alpha}_{\mathcal{U}}$. Therefore,
$\mathrm{MSE}(\sigma^{2},\rho,\mathcal{U},\boldsymbol{\alpha}_{\mathcal{U}})$
$=\frac{1}{|\mathcal{U}|!}\sum_{j=1}^{|\mathcal{U}|!}\mathrm{MSE}(n,\mathcal{U},\sigma^{2},\rho,\Pi_{j}(\boldsymbol{\alpha}))$ (51)
@@
$=\frac{1}{|\mathcal{U}|!}\sum_{j=1}^{|\mathcal{U}|!}\sup_{\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}}\in\mathbb{B}^{d}}f\left(\Pi_{j}(\boldsymbol{\alpha}_{\mathcal{U}}),\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}},\Sigma\right)$ (52)
@@
$\geq\sup_{\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}}\in\mathbb{B}^{d}}\frac{1}{|\mathcal{U}|!}\sum_{j=1}^{|\mathcal{U}|!}f\left(\Pi_{j}(\boldsymbol{\alpha}_{\mathcal{U}}),\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}},\Sigma\right)$ (53)
@@
$\geq\sup_{\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}}\in\mathbb{B}^{d}}f\left(\frac{1}{|\tau|!}\sum_{j=1}^{|\tau|!}(\Pi_{j}(\boldsymbol{\alpha}_{\mathcal{U}})),\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}},\Sigma\right)$ (54)
$=\sup_{\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}}\in\mathbb{B}^{d}}f\left(\mathbf{\tilde{\boldsymbol{\alpha}}_{\mathcal{U}}},\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}},\Sigma\right)$ (55)
where (54) is due to the convexity of $f(\boldsymbol{\alpha}_{\mathcal{U}},\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}},\Sigma)$ in $\boldsymbol{\alpha}_{\mathcal{U}}$, and $\mathbf{\tilde{\boldsymbol{\alpha}}_{\mathcal{U}}}=\left(\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\alpha_{i}\right)\mathbf{1}_{|\mathcal{U}|}$. This implies that for any decoder $\boldsymbol{\alpha}_{\mathcal{U}}$ with $\mathcal{U}$, $\sigma^{2}$ and $\rho$ fixed, there exists a $\mathbf{\tilde{\boldsymbol{\alpha}}}=[\tilde{\alpha}_{j_{1}},\dotsc,\tilde{\alpha}_{j_{|\mathcal{U}|}}]^{T}$ such that $\tilde{\alpha}_{k}=\tilde{\alpha}_{\ell}$, $\forall k,\ell$ satisfying,
$\mathrm{MSE}(\sigma^{2},\rho,\mathcal{U},\boldsymbol{\alpha}_{\mathcal{U}})\geq\mathrm{MSE}(n,\mathcal{U},\sigma^{2},\rho,\mathbf{\tilde{\boldsymbol{\alpha}}_{\mathcal{U}}}).$ (56)
Therefore, for a given $\mathcal{U}\subseteq\mathcal{U}_{all}$, the optimum decoder is of the form $\boldsymbol{\alpha}_{\mathcal{U}}=\alpha\mathbf{1}_{|\mathcal{U}|}$, and (B) can be written as,
$\mathrm{MSE}(n,\mathcal{U},\sigma^{2},\rho,\alpha)$
$=\sup_{\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}}\in\mathbb{B}^{d}}\mathbb{E}\left[\left\|\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\alpha(\mathbf{x}_{i}+\mathbf{Z}_{i})-\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\mathbf{x}_{i}\right\|^{2}\right]$ (57)
$=\sup_{\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}}\in\mathbb{B}^{d}}\frac{1}{|\mathcal{U}|^{2}}\!\sum_{k=1}^{d}\!\left(\!\!(\alpha\!-\!1)^{2}\sum_{i\in\mathcal{U}}x_{i,k}^{2}\right.$
@@
$\quad\left.+(\alpha\!-\!1)^{2}\!\!\sum_{\begin{subarray}{c}i,j\in\mathcal{U}\\ i\neq j\end{subarray}}\!\!x_{i,k}x_{j,k}+\alpha^{2}\sum_{\begin{subarray}{c}i,j\in\mathcal{U}\\ i\neq j\end{subarray}}\rho\sigma^{2}+\alpha^{2}\sigma^{2}|\mathcal{U}|\right)$ (58)
$=\sup_{\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}}\in\mathbb{B}^{d}}\frac{1}{|\mathcal{U}|^{2}}\left(\!\!(\alpha\!-\!1)^{2}\sum_{i\in\mathcal{U}}\|\mathbf{x}_{i}\|^{2}\right.$ (59)
$\left.+(\alpha\!-\!1)^{2}\!\!\sum_{\begin{subarray}{c}i,j\in\mathcal{U}\\ i\neq j\end{subarray}}\mathbf{x}_{i}^{T}\mathbf{x}_{j}+d\alpha^{2}\rho\sigma^{2}|\mathcal{U}|(|\mathcal{U}|-1)+d\alpha^{2}\sigma^{2}|\mathcal{U}|\right)$ (60)
The worst case MSE is resulted when $\mathbf{x}_{i}^{T}\mathbf{x}_{j}=1$, for all $i,j\in\mathcal{U}$, that is, when $\mathbf{x}_{i}=\mathbf{x}_{j}$, $\forall i,j$, and $\mathbf{x}_{i}\in\mathbb{S}^{d-1}$, where $\mathbb{S}^{d-1}$ is the unit sphere in $\mathbb{R}^{d}$. Therefore,
$\mathrm{MSE}(n,\mathcal{U},\sigma^{2},\rho,\alpha)$
$\quad=\frac{1}{|\mathcal{U}|^{2}}\left((\alpha-1)^{2}|\mathcal{U}|+(\alpha-1)^{2}|\mathcal{U}|(|\mathcal{U}|-1)\right.$
$\left.\qquad+d\alpha^{2}\sigma^{2}|\mathcal{U}|(1+\rho(|\mathcal{U}|-1))\right)$ (61)
$\quad=(\alpha-1)^{2}+\frac{d\alpha^{2}\sigma^{2}}{|\mathcal{U}|}(1+\rho(|\mathcal{U}|-1))$ (62)
The optimum decoder for any fixed $\mathcal{U}$, $\sigma^{2}$ and $r$ is computed as,
$\frac{\partial\mathrm{MSE}(n,\mathcal{U},\sigma^{2},\rho,\alpha)}{\partial\alpha}$
$\quad=2(\alpha-1)+\frac{2d\alpha\sigma^{2}}{|\mathcal{U}|}(1+\rho(|\mathcal{U}|-1))=0$ (63)
$\implies\alpha^{*}=\frac{1}{1+\frac{d\sigma^{2}}{|\mathcal{U}|}(1+\rho(|\mathcal{U}|-1))}.$ (64)